fraction of association (?)

What is the fraction of association (?) for the following potassium propionate solutions? Ignore activities. The Ka of propanoic acid is 1.34 × 10-5.
(a) 5.00 × 10-1 M K(C2H5CO2)
(b) 5.00 × 10-2 M K(C2H5CO2)
(c) 5.00 × 10-12 M K(C2H5CO2)
Answer #1

C2H5CO2 + H2O --------> HC2H5CO2 + OH- -----> Kb

With this overall reaction, tthe first thing is calculate Kb:

Kb = Kw/Ka = 1x10-14 / 1.34x10-5 = 7.46x10-10

Now, let's calculate the concentration of the propanoic acid:

C2H5CO2 + H2O --------> HC2H5CO2 + OH-

i) 5x10-1 -----------------------0 0

eq) 5x10-1 - x -------------------x x

7.46x10-10 = x2 / 0.5 - x ----> however Kb is a low value, so we can assume that 0.5 - x = 0.5 so:

x = (7.46x10-10 x 0.5)1/2 = 1.93x10-5 M

Doing the same thing with the other concentration:

x2 = (7.46x10-10 x 0.05)1/2 = 6.11x10-6 M

The α can be calculated like this:

α = [HC2H5CO2] / [C2H5CO2]

α1 = 1.93x10-5 / 0.5 = 3.86x10-5 = α1

α2 = 6.11x10-6 / 0.05 = 1.222x10-4 = α2
Answer #2

C2H5CO2 + H2O --------> HC2H5CO2 + OH- -----> Kb

With this overall reaction, tthe first thing is calculate Kb:

Kb = Kw/Ka = 1x10-14 / 1.34x10-5 = 7.46x10-10

Now, let's calculate the concentration of the propanoic acid:

C2H5CO2 + H2O --------> HC2H5CO2 + OH-

i) 5x10-1 -----------------------0 0

eq) 5x10-1 - x -------------------x x

7.46x10-10 = x2 / 0.5 - x ----> however Kb is a low value, so we can assume that 0.5 - x = 0.5 so:

x = (7.46x10-10 x 0.5)1/2 = 1.93x10-5 M

Doing the same thing with the other concentration:

x2 = (7.46x10-10 x 0.05)1/2 = 6.11x10-6 M

The α can be calculated like this:

α = [HC2H5CO2] / [C2H5CO2]

α1 = 1.93x10-5 / 0.5 = 3.86x10-5 = α1

α2 = 6.11x10-6 / 0.05 = 1.222x10-4 = α2